OpenStudy (anonymous):

Which system has no solutions? Which system has infinite solutions? Which system has (1, 3) as the solution? Which system has (-8, 0) as the solution? Which system has (-2, -2) as the solution?

OpenStudy (anonymous):

4x+6y=-12 2x+3y=9 x-y=-2 3x-y=0

OpenStudy (anonymous):

A system with infinite solutions are the same line. For a point (x,y) to be a solution, substituted the x and y into the eqation. A system with no solutions are parallel.

OpenStudy (anonymous):

x+y=-4 -3x+2y=2 x+4=-1 4x+4y=-4

OpenStudy (anonymous):

x-4y=-8 2x-3y=-16

OpenStudy (anonymous):

help please will fan and medal

OpenStudy (anonymous):

4x+6y=-12 2x+3y=9 x-y=-2 3x-y=0 x+y=-4 -3x+2y=2 x+4=-1 4x+4y= -4x-4y=-8 2x-3y=-16

OpenStudy (acxbox22):

for a solution of (1,3) you can graph any 2 lines that intersect at (1,3) and write down the equations for the 2 lines as the system use this process for the next 2

OpenStudy (anonymous):

CANT U JUST GIVE ME THE ANSWERS

OpenStudy (acxbox22):

sorry, it is against the code of conduct

OpenStudy (acxbox22):

i know i sound dorky but the mods ban you

OpenStudy (anonymous):

please

OpenStudy (acxbox22):

and i need some hw help myself so i cant afford that

OpenStudy (anonymous):

ill help you if you help me

OpenStudy (acxbox22):

do you know exponential functions?